# Quantifying Holistic Review: A Multi-Modal Approach to College Admissions Prediction

### v1

@@ III-B1 Step 1: Hybrid Weighting Strategy
$X_{\text{scaled}}=\text{StandardScaler}(X)$
@@
$\text{PC}_{1},\text{PC}_{2}=\text{PCA}(X_{\text{scaled}})$
@@
$w_{\text{PCA}}^{\text{raw}}=-(\alpha_{\text{PCA}}\cdot\text{PC}_{1}+\beta_{\text{PCA}}\cdot\text{PC}_{2})$
@@
$w_{\text{PCA}}=\frac{w_{\text{PCA}}^{\text{raw}}}{\sum_{i}w_{\text{PCA},i}^{\text{raw}}}$
@@
$w_{\text{fused}}=\alpha_{\text{fusion}}\cdot w_{\text{PCA}}+(1-\alpha_{\text{fusion}})\cdot w_{\text{manual}}$
@@ III-C2 Prompt Alignment and Adjustment
$\text{EQI}_{\text{final}}=\text{EQI}_{\text{raw}}\times\left(\lambda+(1-\lambda)\cdot\frac{1}{1+e^{-k(s_{\text{align}}-x_{0})}}\right)$
@@ III-D EIS Module
[…]
@@
[…]
@@ III-E CAPS Score Fusion Module
$w_{i}=\alpha\cdot w_{\text{log},i}+\beta\cdot w_{\text{xgb},i}+\gamma\cdot w_{\text{exp},i},\quad\text{where }\alpha{+}\beta{+}\gamma=1$
We use $(\alpha,\beta,\gamma)=(0.3,0.3,0.4)$ by default. The final weights are then normalized so that $\sum_{i}w_{i}=1$.
@@
$\text{CAPS}_{\text{raw}}=\sum_{i}w_{i}\cdot x_{i}$
@@
$\text{CAPS}_{\text{final}}=\min(100,\text{CAPS}_{\text{raw}}\times 100+\text{bonus})$
@@ IV-1 Model Setup
[…]
@@
$\textbf{MSE}=0.0316,\quad\textbf{R}^{2}=0.7999$
@@
Figure 7: SHAP plot showing key feature contributions in EQI.
@@ IV-A Model Evaluation and Robustness
The macro-averaged F1 score was 0.69, and the weighted F1 score reached 0.74, indicating strong linear separability for most tiers, albeit with some underfitting in higher tiers due to limited sample diversity.
@@
V Conclusion

### v2

@@ III-B1 Step 1: Hybrid Weighting Strategy
$X_{\text{scaled}}=\text{StandardScaler}(X)$ (2)
@@
$\text{PC}_{1},\text{PC}_{2}=\text{PCA}(X_{\text{scaled}})$ (3)
@@
$w_{\text{PCA}}^{\text{raw}}=-(\alpha_{\text{PCA}}\cdot\text{PC}_{1}+\beta_{\text{PCA}}\cdot\text{PC}_{2})$ (4)
@@
$w_{\text{PCA}}=\frac{w_{\text{PCA}}^{\text{raw}}}{\sum_{i}w_{\text{PCA},i}^{\text{raw}}}$ (5)
@@
$w_{\text{fused}}=\alpha_{\text{fusion}}\cdot w_{\text{PCA}}+(1-\alpha_{\text{fusion}})\cdot w_{\text{manual}}$ (6)
@@ III-C2 Prompt Alignment and Adjustment
$\text{EQI}_{\text{final}}=\text{EQI}_{\text{raw}}\times\left(\lambda+(1-\lambda)\cdot\frac{1}{1+e^{-k(s_{\text{align}}-x_{0})}}\right)$ (10)
@@ III-D EIS Module
$\text{EIS}_{\text{activity}}=\gamma\cdot\text{GPT\_Score}+(1-\gamma)\cdot\text{Tier\_Score}$ (12)
@@
$\text{EIS}_{\text{final}}=\text{Avg\_EIS}\times(0.85+0.15\cdot\text{Coherence\_Score})$ (13)
@@ III-E CAPS Score Fusion Module
$w_{i}=\alpha\cdot w_{\text{log},i}+\beta\cdot w_{\text{xgb},i}+\gamma\cdot w_{\text{exp},i}$ (14)
We use $(\alpha,\beta,\gamma)=(0.3,0.3,0.4)$ by default. This specific weighting was determined through preliminary experimentation, which indicated that a slight emphasis on expert priors ($\gamma=0.4$) stabilized the model against noise from the data-driven weights, while still giving substantial influence to both the linear ($\alpha=0.3$) and non-linear ($\beta=0.3$) patterns. This balanced approach ensures the model is robust and well-aligned with established admissions heuristics. The final weights are then normalized so that $\sum_{i}w_{i}=1$.
@@
$\text{CAPS}_{\text{raw}}=\sum_{i}w_{i}\cdot x_{i}$ (15)
@@
$\text{CAPS}_{\text{final}}=\min(100,\text{CAPS}_{\text{raw}}\times 100+\text{bonus})$ (16)
@@ IV-1 Model Setup
We performed an 80/20 train/test split with a fixed random seed ($\texttt{random\_state}=42$). Model selection was performed via exhaustive grid search with 3-fold cross-validation using the following hyperparameter space:
@@
$\textbf{MSE}=0.0316,\quad\textbf{R}^{2}=0.7999$ (18)
@@
Figure 7: SHAP Summary Plot for the EQI XGBoost Model. This plot displays the top 15 features ranked by their impact on the model’s prediction of the Essay Quality Index (EQI). Each point represents an essay in the dataset. The feature’s position on the y-axis indicates its importance. The x-axis shows the SHAP value, where positive values increase the predicted EQI score and negative values decrease it. The color indicates the feature’s value, with red being high and blue being low. For instance, high values for EssayLanguageScore (red dots) have high positive SHAP values, showing a strong, positive contribution to the final EQI score.
@@ IV-A Model Evaluation and Robustness
a) Multinomial Logistic Regression.: A multinomial logistic regression model was trained on Z-scored features using an $80/20$ train-test split with a fixed random seed (random_state 42). The model achieved an overall accuracy of 75% on the test set. Detailed precision, recall, and F1 scores across the five tiers are shown below:
  • Tier 0-2: Precision/Recall/F1 = 1.00
  • Tier 3: Precision =0.40=0.40, Recall = 0.50, $F1=0.44$
  • Tier 4: Precision =0.00,=0.00, $Recall=0.00$, $F1=0.00$
The model’s complete failure to classify Tier 4 applicants is a direct consequence of significant class imbalance within our synthetic dataset, where this tier was sparsely represented. This limitation highlights a common challenge in modeling rare outcomes in admissions data.
The macro-averaged F1 score was 0.69, and the weighted F1 score reached 0.74, indicating strong linear separability for most tiers. However, the poor performance on Tier 4 suggests that while the features are informative, the model requires mitigation strategies for imbalanced classes. Future work could address this by employing techniques such as oversampling the minority class (e.g., SMOTE), applying class weights during model training, or collecting a more balanced dataset.
@@
V Limitations
While the CAPS framework demonstrates strong potential, we acknowledge several limitations that should be addressed in future work.
First, the primary limitation is the use of a synthetic dataset. Although constructed to be realistic by reflecting known distributions and correlations in admissions data, it cannot capture the full complexity, noise, and nuanced interdependencies of a real-world applicant pool. Consequently, the model’s performance metrics should be interpreted as a proof-of-concept rather than a direct measure of real-world efficacy.
Second, the generalizability of our findings is constrained. The component weightings and model performance are specific to the characteristics of our dataset. Different types of institutions (e.g., large public universities vs. small liberal arts colleges) have distinct evaluation criteria, and the CAPS framework would require retraining and validation on their specific historical data to be applicable.
Finally, while we use LLMs to score qualitative components like essays and extracurriculars, these models can have inherent biases. Ensuring fairness and mitigating potential biases from LLM-generated scores is a critical area for further research before deploying such a system in a high-stakes environment.
VI Conclusion